OpenStudy (anonymous):

In a certain year the amount of a of garbage in pounds produced after t days by a avg house hold is given by A=9.5t A graph the equation for t>0 B how many days di it take the avg house hold to produce 250 lb

OpenStudy (ivancsc1996):

A. The graph is a straight line starting at (0,0) and increasing 9.5 in the y-axis, by every step in the x-axis. (Label A the y-axis and t the x-axis) B. t=A/9.5=250/9.5